Company A = mean 1300
= Standard deviation 82
.= Sample number 100
Company B= mean 1248
.... = Standard deviation 93
= Sample number 100
Which brand of bulbs are you going to buy if you wish to take risk at 5%
1
Expert's answer
2021-07-29T14:24:32-0400

Company A

"n_1=100 \\

\bar{x}_1 = 1300 \\

s_1=80"

Company B

"n_2=100 \\

\bar{x}_2=1248 \\

s_2=93 \\

H_0: \mu_1= \mu_2 \\

H_1: \mu_1 ≠ \mu_2"

Test-statistic: two-sample t test

"t = \frac{\bar{x_1} -\bar{x_2}}{s_p \sqrt{ \frac{1}{n_1} + \frac{1}{n_2} }} \\

s^2_p= \frac{(n_1-1)s^2_1 + (n_2-1)s^2_2}{n_1+n_2-2} \\

s^2_p = \frac{(100-1) \times (80)^2 + (100-1) \times (93)^2}{100+100-2} \\

= \frac{633600+856251}{198} \\

= 7524.5 \\

s_p= \sqrt{7524.5}=86.74 \\

t = \frac{1300 -1248}{86.74 \sqrt{ \frac{1}{100} + \frac{1}{100} }} \\

= \frac{52}{12.26} \\

= 4.24"

Tabulated value of t at 5% level of significance and d.f.= 100 +100 – 2 = 198 is 1.972.

Since the calculated value of t is greater than the tabulated value of t at 5% level of significance, H0 is rejected. There is a significant difference in the mean life of bulbs of the two companies.

I am going to buy bulbs of brand A.